# Proposal for simulating quantum spin models with Dzyaloshinskii-Moriya interaction using Rydberg atoms, and construction of asymptotic quantum many-body scar states

### v4

@@ V.2 Open boundary case
Here, we defined $\hat{S}_{0}^{z}=\hat{S}_{M+1}^{z}=-1/2$. The last term in Eq. (30a) is necessary to define the soliton number operator that commutes with Eq. (30a). In Eq. (30b), we add the auxiliary sites $j=0$, $M+1$ and restrict the states to those with the down spins at the two sites. This restriction is equivalent to applying a downward magnetic field with infinite strength to the spins at the auxiliary sites.
@@ V.3 Level spacing statistics
Figure 6: $r$-value as a function of the magnetic field. The green solid and black dashed lines represent the $r$-value of the Wigner-Dyson (Gaussian orthogonal ensemble) and Poisson distribution, respectively. (a) DH model without the edge magnetic field for $M=16$. (b) DH model with the edge magnetic field for $M=18$ in $N_{\rm sol}^{\rm OBC}=5$ sector.
Figure 7: $r$-value as a function of the magnetic field in the low-field regime. The green solid and black dashed lines represent the $r$-value of the Wigner-Dyson (Gaussian orthogonal ensemble) and Poisson distribution, respectively. (a) DH model without the edge magnetic field for $M=16$. (b) DH model with the edge magnetic field for $M=18$ in $N_{\rm sol}^{\rm OBC}=5$ sector.
@@
Figure 6 (a) shows the results of the magnetic-field dependence of the $r$-value of the DH model without an edge magnetic field. In this case, $\hat{C}$ is conserved, but $\hat{N}_{\rm sol}^{\rm OBC}$ is not. The $r$-value is consistent with the Wigner-Dyson distribution (Gaussian orthogonal ensemble) around $h^{x}\lesssim D$ ($r\simeq 0.53$). In the high-field regime, the $r$-value approaches the Poisson value ($r\simeq 0.38$), due to the emergent conservation law of $\hat{S}_{\rm tot}^{z}$ in the high-field regime. We also plot the $r$-value of the DH model with the edge magnetic field in Fig. 6 (b), where we calculate the largest symmetry sector. For $M=18$, e.g,, $N_{\rm sol}^{\rm OBC}=5$ is the largest.
Figures 7 (a) and (b) represent the $r$-value in the low-field regime. We can see the deviation of the $r$-value from the Wigner-Dyson value around $h^{x}\sim 0.01D$. This behavior results from the integrability at $h^{x}=0$; the DH model is equivalent to the XY model at this point in the absence of the edge magnetic field. The integrability of the XY model with the edge magnetic field terms was discussed in Refs. Bilstein1999 ; Bilstein2000 ; Bilstein2000_2 . Regardless of the presence/absence of the edge magnetic field, the $r$-value does not approach the Poisson value. We may attribute this discrepancy to the additional emergent symmetry at $h^{x}=0$ discussed in Refs. Bilstein1999 ; Bilstein2000 ; Bilstein2000_2 .
@@ V.4 Entanglement entropy and off-diagonal long-range order
In this subsection, we discuss the half-chain entanglement entropy of the system under the open boundary condition. Figure 8 (a) shows the half-chain von Neumann EE as a function of the eigenenergy of the system. We find a QMBS state with a high energy eigenstate and low EE. One can derive the analytical expression of the EE of the QMBS state (see Appendix E for details). We show the size dependence of the EE in Fig. 8 (b). The EE obeys the sub-volume law $\sim\ln M$.
@@ VI Asymptotic quantum many-body scar states of the DH model
Figure 10: System size dependence of the half-chain EE for the states $\left|AS_{n}\right\rangle$ for several $n$.
@@
Figure 11: The energy variance of the states $\left|AS_{n}\right\rangle$ for several $n$ as a function of the system size $M$. “Leading" in the legend means $\Delta E^{2}/D^{2}=\pi^{2}/(4M^{2})$.
Finally, we discuss the dynamics after quench from the state $\left|AS_{n}\right\rangle$. Figures 12 (a), (b), and (c) show the fidelity $|\left\langle AS_{n}\right|e^{-i\hat{H}_{\rm DH}^{\rm OBC}t/\hbar}\left|AS_{n}\right\rangle|^{2}$ for several system sizes. We use the TEBD method to obtain these results. The discrete time step is set to $\Delta t=0.01\hbar/D$, and the bond dimension is taken to be $\chi=200\sim 310$ so that the truncation error is less than $10^{-10}$. At early times, the fidelity can be approximated as $|\left\langle AS_{n}\right|\exp(-i\hat{H}_{\rm DH}^{\rm OBC}t/\hbar)\left|AS_{n}\right\rangle|^{2}\simeq 1-\Delta E^{2}t^{2}/\hbar^{2}$, which implies that the smaller energy variance yields a longer relaxation time. As expected, we see the system size dependence of the fidelity even for $n>1$. We also plot the rescaled fidelity in Figs. 12(d), (e), and (f). We can find a good data collapse by rescaling $t/M$. These results all suggest that the state (41) is an AQMBS state in the DH model.
Figure 12: Time evolution of the fidelity with the initial condition (a) $\left|AS_{1}\right\rangle$, (b) $\left|AS_{5}\right\rangle$, (c) $\left|AS_{10}\right\rangle$. (d), (e), and (f): Rescaled fidelity of (a), (b), and (c), respectively.
@@ VII Summary
In this paper, we proposed a new method for performing quantum simulations of the $S=1/2$ quantum spin model with the DMI using Rydberg atoms. Our approach is based on the two-photon Raman scheme and unitary transformation, which is feasible with current experimental techniques. One advantage of the Rydberg atom quantum simulator is that it realizes quantum spin models with high tunability. In our scheme, we can easily tune the ratio between the magnitude of the exchange interaction and the DMI in a wide range. This tunability allows us to implement the DH model, which has only the DMI and Zeeman energy. Another advantage is that the heating problem is insignificant compared to the SOC case because the typical timescale is much shorter than the neutral atom case due to the strong interaction between the Rydberg atoms.
@@
Our work opens a new avenue for the design of quantum chiral magnets. There are several future directions. One is to extend our scheme to a more complex form of the DMI, which would enable experimental studies of skyrmions in the quantum regime Ochoa2019 . Another direction is an extension to higher spin systems, allowing us to explore the spin parity effect in the DH model.
@@ Appendix A Derivation of the XXZ Hamiltonian
In this appendix, we discuss the interaction between Rydberg states $|n_{1}S_{1/2},m_{J}=+1/2\rangle\equiv\left|\downarrow\right\rangle$ and $|n_{2}S_{1/2},m_{J}=+1/2\rangle\equiv\left|\uparrow\right\rangle$ with energy $E_{n_{1}S_{1/2}}$ and $E_{n_{2}S_{1/2}}$. Since these two states have the same parity, they are not coupled directly via the dipole-dipole interaction. If the atomic distance is large enough, the dipole-dipole interaction can be treated as a perturbation, and the leading term is given by the second-order term Bijnen_thesis ; Whitlock2017 . For simplicity, we consider only the two Rydberg atoms separated by the distance $R$. The Hamiltonian is given by Weber2017
@@ Appendix C Relation between the spin-laboratory and spin-rotating frames
Table 1: List of the correspondence between the spin-laboratory and spin-rotating frames. $\left|\uparrow_{j}\right\rangle$ and $\left|\downarrow_{j}\right\rangle$ represent the eigenstates of $\hat{S}_{j}^{z}$ with eigenvalue $+1/2$ and $-1/2$, respectively.
@@ Appendix D Extension to two-dimensional systems
Figure 13: Schematic of the setup for the experimental realization of the DMI in a two-dimensional square lattice. The filled black circles represent the position of the Rydberg atoms. The magnetic field is applied along the $z$ axis.
@@ Appendix E Analytical expression of the von Neumann entanglement entropy of the quantum many-body scar states
where $k$ represents the number of up spins in the subsystem $A$, and $K\equiv\min({n,1+\lfloor M/4\rfloor})$. The element of the matrix $M^{n,k}_{\bm{m}_{A},\bm{m}_{B}}$ takes the value $1/\sqrt{N(M,n)}$ [see Eq. (36)] if $\left|\bm{m}_{A}\right\rangle$ contains $k$-up spins and $\left|\bm{m}_{B}\right\rangle$ contains $(n-k)$-up spins and the constraint of the spin configuration is satisfied, otherwise zero.
@@
$D_{1,k}$ $\equiv N(M/2-1,k),$ (80)
$D_{2,k}$ $\equiv N(M/2-2,k-1),$ (81)
$l_{1,k}$ $\equiv N(M/2,n-k),$ (82)
$l_{2,k}$ $\equiv N(M/2-1,n-k),$ (83)
@@
$\frac{1}{N(M,n)}\begin{bmatrix}l_{1,k}D_{1,k}&l_{2,k}D_{2,k}\\ l_{2,k}D_{1,k}&l_{2,k}D_{2,k}\end{bmatrix}.$ (84)
@@
$\lambda_{k,\pm}$ $=\frac{1}{2N(M,n)}\left[l_{1,k}D_{1,k}+l_{2,k}D_{2,k}\pm\sqrt{(l_{1,k}D_{1,k}-l_{2,k}D_{2,k})^{2}+4D_{1,k}D_{2,k}l_{2,k}^{2}}\right].$ (85)
@@ Appendix F Edge state in the Krylov space
$\hat{H}\left|u_{j}\right\rangle$ $=\alpha_{j}\left|u_{j}\right\rangle+\beta_{j-1}\left|u_{j-1}\right\rangle+\beta_{j}\left|u_{j+1}\right\rangle,\quad(j=1,2,\cdots,N_{\rm K})$ (87)
@@
Figure 14: (a) Krylov coefficient $\beta_{i}$ for $M=16$ and $h^{x}=0.1D$. (b) Magnified view of (a).
From the above results, we can write down the matrix representation of the DH Hamiltonian as
@@
We can regard this matrix as a one-dimensional tight-binding Hamiltonian with nonuniform nearest-neighbor hopping amplitude $\beta_{j}$. We can easily find that the unitary matrix $\Gamma={\rm diag}(1,-1,1,-1,\ldots)$ satisfies the relation $\{\Gamma,H_{\rm DH}\}=0$. Because all the matrix elements of the matrix $H_{\rm DH}$ are real, we find that this Hamiltonian belongs to the symmetry class BDI in the Altland-Zirnbauer classification Altland1997 . According to the periodic table of topological insulators and superconductors, edge states are allowed in one-dimensional class BDI systems Schnyder2008 .

### v5

@@ V.2 Open boundary case
Here, we defined $\hat{S}_{0}^{z}=\hat{S}_{M+1}^{z}=-\frac{1}{2}$. The last term in Eq. (30a) is necessary to define the soliton number operator that commutes with Eq. (30a). In Eq. (30b), we add the auxiliary sites $j=0$, $M+1$ and restrict the states to those with the down spins at the two sites. This restriction is equivalent to applying a downward magnetic field with infinite strength to the spins at the auxiliary sites.
@@ V.3 Level spacing statistics
Figure 6: $r$ value as a function of the magnetic field. The green solid and black dashed lines represent the $r$ value of the Wigner-Dyson (Gaussian orthogonal ensemble) and Poisson distribution, respectively. (a) DH model without the edge magnetic field for $M=16$. (b) DH model with the edge magnetic field for $M=18$ in $N_{\rm sol}^{\rm OBC}=5$ sector.
Figure 7: $r$ value as a function of the magnetic field in the low-field regime. The green solid and black dashed lines represent the $r$ value of the Wigner-Dyson (Gaussian orthogonal ensemble) and Poisson distribution, respectively. (a) DH model without the edge magnetic field for $M=16$. (b) DH model with the edge magnetic field for $M=18$ in $N_{\rm sol}^{\rm OBC}=5$ sector.
@@
Figure 6 (a) shows the results of the magnetic-field dependence of the $r$ value of the DH model without an edge magnetic field. In this case, $\hat{C}$ is conserved, but $\hat{N}_{\rm sol}^{\rm OBC}$ is not. The $r$ value is consistent with the Wigner-Dyson distribution (Gaussian orthogonal ensemble) around $h^{x}\lesssim D$ ($r\simeq 0.53$). In the high-field regime, the $r$ value approaches the Poisson value ($r\simeq 0.38$), due to the emergent conservation law of $\hat{S}_{\rm tot}^{z}$ in the high-field regime. We also plot the $r$ value of the DH model with the edge magnetic field in Fig. 6 (b), where we calculate the largest symmetry sector. For $M=18$, e.g., $N_{\rm sol}^{\rm OBC}=5$ is the largest.
Figures 7 (a) and (b) represent the $r$ value in the low-field regime. We can see the deviation of the $r$ value from the Wigner-Dyson value around $h^{x}\sim 0.01D$. This behavior results from the integrability at $h^{x}=0$; the DH model is equivalent to the XY model at this point in the absence of the edge magnetic field. The integrability of the XY model with the edge magnetic field terms was discussed in Refs. Bilstein1999 ; Bilstein2000 ; Bilstein2000_2 . Regardless of the presence or absence of the edge magnetic field, the $r$ value does not approach the Poisson value. We may attribute this discrepancy to the additional emergent symmetry at $h^{x}=0$ discussed in Refs. Bilstein1999 ; Bilstein2000 ; Bilstein2000_2 .
@@ V.4 Entanglement entropy and off-diagonal long-range order
In this subsection, we discuss the half-chain entanglement entropy of the system under the open boundary condition. Figure 8 (a) shows the half-chain von Neumann EE as a function of the eigenenergy of the system. We find a QMBS state with a high-energy eigenstate and low EE. One can derive the analytical expression of the EE of the QMBS state (see Appendix E for details). We show the size dependence of the EE in Fig. 8 (b). The EE obeys the sub-volume law $\sim\ln{M}$.
@@ VI Asymptotic quantum many-body scar states of the DH model
Figure 10: System-size dependence of the half-chain EE for the states $\left|AS_{n}\right\rangle$ for several $n$.
@@
Figure 11: The energy variance of the states $\left|AS_{n}\right\rangle$ for several $n$ as a function of the system size $M$. “Leading” in the legend means $\Delta E^{2}/D^{2}=\pi^{2}/(4M^{2})$.
Finally, we discuss the dynamics after quench from the state $\left|AS_{n}\right\rangle$. Figures 12(a)-12(c) show the fidelity $|\left\langle AS_{n}\right|e^{-i\hat{H}_{\rm DH}^{\rm OBC}t/\hbar}\left|AS_{n}\right\rangle|^{2}$ for several system sizes. We use the TEBD method to obtain these results. The discrete time step is set to $\Delta t=0.01\hbar/D$, and the bond dimension is taken to be $\chi=200\sim 310$ so that the truncation error is less than $10^{-10}$. At early times, the fidelity can be approximated as $|\left\langle AS_{n}\right|\exp(-i\hat{H}_{\rm DH}^{\rm OBC}t/\hbar)\left|AS_{n}\right\rangle|^{2}\simeq 1-\Delta E^{2}t^{2}/\hbar^{2}$, which implies that the smaller energy variance yields a longer relaxation time. As expected, we see the system size dependence of the fidelity even for $n>1$. We also plot the rescaled fidelity in Figs. 12(d)-12(f). We can find a good data collapse by rescaling $t/M$. These results all suggest that the state (41) is an AQMBS state in the DH model.
Figure 12: Time evolution of the fidelity with the initial condition (a) $\left|AS_{1}\right\rangle$, (b) $\left|AS_{5}\right\rangle$, (c) $\left|AS_{10}\right\rangle$. (d)–(f): Rescaled fidelity of (a), (b), and (c), respectively.
@@ VII Summary
In this paper, we proposed a new method for performing quantum simulations of the $S=\frac{1}{2}$ quantum spin model with the DMI using Rydberg atoms. Our approach is based on the two-photon Raman scheme and unitary transformation, which is feasible with current experimental techniques. One advantage of the Rydberg atom quantum simulator is that it realizes quantum spin models with high tunability. In our scheme, we can easily tune the ratio between the magnitude of the exchange interaction and the DMI in a wide range. This tunability allows us to implement the DH model, which has only the DMI and Zeeman energy. Another advantage is that the heating problem is insignificant compared to the SOC case because the typical timescale is much shorter than the neutral atom case due to the strong interaction between the Rydberg atoms.
@@
Our work opens a new avenue for the design of quantum chiral magnets. There are several future directions. One is to extend our scheme to a more complex form of the DMI, which would enable experimental studies of skyrmions in the quantum regime Ochoa2019 . Another direction is an extension to higher spin systems, allowing us to explore the spin-parity effect in the DH model.
@@ Appendix A Derivation of the XXZ Hamiltonian
In this appendix, we discuss the interaction between Rydberg states $|n_{1}S_{1/2},m_{J}=+\frac{1}{2}\rangle\equiv\left|\downarrow\right\rangle$ and $|n_{2}S_{1/2},m_{J}=+\frac{1}{2}\rangle\equiv\left|\uparrow\right\rangle$ with energy $E_{n_{1}S_{1/2}}$ and $E_{n_{2}S_{1/2}}$. Since these two states have the same parity, they are not coupled directly via the dipole-dipole interaction. If the atomic distance is large enough, the dipole-dipole interaction can be treated as a perturbation, and the leading term is given by the second-order term Bijnen_thesis ; Whitlock2017 . For simplicity, we consider only the two Rydberg atoms separated by the distance $R$. The Hamiltonian is given by Weber2017
@@ Appendix C Relation between the spin-laboratory and spin-rotating frames
Table 1: List of the correspondence between the spin-laboratory and spin-rotating frames. $\left|\uparrow_{j}\right\rangle$ and $\left|\downarrow_{j}\right\rangle$ represent the eigenstates of $\hat{S}_{j}^{z}$ with eigenvalue $+\frac{1}{2}$ and $-\frac{1}{2}$, respectively.
@@ Appendix D Extension to two-dimensional systems
Figure 13: Schematics of the setup for the experimental realization of the DMI in a two-dimensional square lattice. The filled black circles represent the position of the Rydberg atoms. The magnetic field is applied along the $z$ axis.
@@ Appendix E Analytical expression of the von Neumann entanglement entropy of the quantum many-body scar states
where $k$ represents the number of up spins in the subsystem $A$, and $K\equiv\min({n,1+\lfloor M/4\rfloor})$. The element of the matrix $M^{n,k}_{\bm{m}_{A},\bm{m}_{B}}$ takes the value $1/\sqrt{\mathcal{N}(M,n)}$ [see Eq. (36)] if $\left|\bm{m}_{A}\right\rangle$ contains $k$-up spins and $\left|\bm{m}_{B}\right\rangle$ contains $(n-k)$-up spins and the constraint of the spin configuration is satisfied, otherwise zero.
@@
$D_{1,k}$ $\equiv\mathcal{N}(M/2-1,k),$ (80)
$D_{2,k}$ $\equiv\mathcal{N}(M/2-2,k-1),$ (81)
$l_{1,k}$ $\equiv\mathcal{N}(M/2,n-k),$ (82)
$l_{2,k}$ $\equiv\mathcal{N}(M/2-1,n-k),$ (83)
@@
$\frac{1}{\mathcal{N}(M,n)}\begin{bmatrix}l_{1,k}D_{1,k}&l_{2,k}D_{2,k}\\ l_{2,k}D_{1,k}&l_{2,k}D_{2,k}\end{bmatrix}.$ (84)
@@
$\lambda_{k,\pm}$ $=\frac{1}{2\mathcal{N}(M,n)}\left[l_{1,k}D_{1,k}+l_{2,k}D_{2,k}\pm\sqrt{(l_{1,k}D_{1,k}-l_{2,k}D_{2,k})^{2}+4D_{1,k}D_{2,k}l_{2,k}^{2}}\right].$ (85)
@@ Appendix F Edge state in the Krylov space
$\hat{H}\left|u_{j}\right\rangle$ $=\alpha_{j}\left|u_{j}\right\rangle+\beta_{j-1}\left|u_{j-1}\right\rangle+\beta_{j}\left|u_{j+1}\right\rangle,\quad(j=1,2,\cdots,N_{\rm K}),$ (87)
@@
Figure 14: (a) Krylov coefficients $\beta_{i}$ for $M=16$ and $h^{x}=0.1D$. (b) Magnified view of (a).
From the above results, we can write the matrix representation of the DH Hamiltonian as
@@
We can regard this matrix as a one-dimensional tight-binding Hamiltonian with nonuniform nearest-neighbor hopping amplitude $\beta_{j}$ (see Fig. 14). We can easily find that the unitary matrix $\Gamma={\rm diag}(1,-1,1,-1,\ldots)$ satisfies the relation $\{\Gamma,H_{\rm DH}\}=0$. Because all the matrix elements of the matrix $H_{\rm DH}$ are real, we find that this Hamiltonian belongs to the symmetry class BDI in the Altland-Zirnbauer classification Altland1997 . According to the periodic table of topological insulators and superconductors, edge states are allowed in one-dimensional class BDI systems Schnyder2008 .